OpenStudy (anonymous):

find the value of 3x (over) (sqrt)x3

OpenStudy (mimi_x3):

is it: \[\frac{3x}{\sqrt{3x}} ?\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\frac{3x}{\sqrt{x^3}}=\frac{3x*\sqrt{x^3}}{\sqrt{x^3}*\sqrt{x^3}}\]\[=\frac{3x*\sqrt{x*x^2}}{x^3}\]\[=\frac{3x^2\sqrt{x}}{x^3}\]\[=\frac{3\sqrt{x}}{x}\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

wait is it \[\frac{3x}{\sqrt{x^3}}\] or \[\frac{3x}{\sqrt{3x}}\] ?

OpenStudy (mimi_x3):

if its : \[\frac{3x}{\sqrt{3x}} \] then.. \[\frac{3x*\sqrt{3x}}{\sqrt{3x}*\sqrt{3x}} =>\frac{3x*\sqrt{3x}}{3x} =\sqrt{3x} \]